my name is jasmine i attend bedford high school i always have problems with fractions i need your help whats
1 2/3 + 3 1/4

Answer by Alan3354(21587)

1 2/3 + 3 1/4
Convert to fractions first.
1 2/3 = 5/3
3 1/4 = 13/4
-----------
=(5/3) + (13/4)
Now find a common denominator. In class, they fret about the LCD, least common denominator, but anyone will do. What ALWAYS works is to multiply the 2 DENs.
3*4 = 12
In this case, it happens to be the LCD, but it isn't always, but it always works.
Now, find what 5/3 is with the DEN of 12. You multiply NUM and DEN by 4.
5/3 = 20/12
Do the same with 13/4, but multiply by 3 to get the DEN of 12.
13/4 = 39/12
--------
Then add them.
(20/12) + (39/12) = 59/12. That's it.
You can make it a mixed fraction. 59/12 = 4 11/12, but I see no advantage to that.
BTW, I have a 5 yr old niece named Jasmine.